Against the wind a commercial airline in South America flew 340 miles in 2.5 hours With a tailwind the return trip took 2 hours. What was the speed of the airplane in still air? What was the speed of the wind? thank you for answering!!
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 2.5 hours
with wind 2 hours

Distance with wind 340 miles distance against wind 340
t=d/r against wind (x-y)
340.00 / ( x - y )= 2.50

2.50 x - -2.50 y = 340.00 ....................1
with wind (x+y)
340.00 / ( x + y )= 2.00
2.00 ( x + y ) = 340.00
2.00 x + 2.00 y = 340.00 ...............2
Multiply (1) by 2.00
Multiply (2) by 2.50
we get 2.00
5.00 x + -5.00 y = 680.00
5.00 x + 5.00 y = 850.00
10.00 x = 1530.00
/ 10.00
x = 153 mph

plug value of x in (1) y
2.5 x -2.5 y = 340
382.5 -2.5 -382.5 = 340
-2.5 y = 340
-2.5 y = -42.5 mph
y = 17
Plane 153 mph
wind 17 mph